OpenStudy (anonymous):

How much money will I make if I work for 6 months, 12 dollars an hour, and 5 days a week

OpenStudy (anonymous):

how many hours r u planning on working? And do u have a money goal?

OpenStudy (anonymous):

for example if you wanted to work 5 hours on a five day week for six months then your equation would look like this 5*12=60 60*5= $300 and since there are about 26 weeks in 6 months 60*26= $1560 this is only assuming you are working for 5 hours. If there is a different amount tell me and we can work it out together I hoped it helped :)

OpenStudy (anonymous):

@Shadowdragon you made a mistake when 5 hour in a day, and 5 day in a week then for 26 weeks 26*5=130 days 130*5=650 hours and the 650*12= 7800$

OpenStudy (anonymous):

Oh sorry my bad this is what you get for rushing